# Relational Self-supervised Distillation with Compact Descriptors for Image Copy Detection

### v1

@@ 5.4. Ablation Study
Effectiveness of Each Loss Term. To further assess the effectiveness of RDCD, we investigate the impact of different loss components on its performance, specifically, RSD loss and HN loss as shown in Table 6. In this section, we present an additional metric, the rank preserving ratio (RPR), which is the rank divided by the descriptor dimension. A low RPR indicates dimensional collapse, and RPR value of 1 indicates that the descriptor makes full use of all dimensions. All models are trained on the DISC21 dataset with 100 epochs and SSCD-RN50-w/mixup is employed as the teacher model. Our result shows that with the addition of RSD loss and HN loss, the performance improves substantially, which indicates that RSD and HN loss can benefit the performance of RDCD.
@@
Table 6. Ablation study on the loss terms in RDCD. Descriptor sizes of 64, 128, and 256 were examined, and a consistent MLP configuration was maintained across all experiments for a fair comparison.
@@
Influence of $\lambda_{rel}$ and $\lambda_{hn}$. $\lambda_{rel}$ and $\lambda_{hn}$ are the weights of RSD loss and HN loss, respectively. First, we analyze the influence of contrastive loss and RSD loss (Table 7). As shown in the third row, a contrastive loss of 1 and RSD loss of 10 achieves the best performance, which indicates that knowledge distillation is essential for the model with better performance. The HN loss is also varied with the contrastive loss and RSD loss fixed at 1 and 10, respectively, showing that the best performance occurs when the HN loss is 5. However, for larger values, the performance drops to 0, (i.e. collapses) because the network focuses more on HN loss rather than contrastive and distillation loss. We further calculate RPR of each experiment. In the first section, where only contrastive loss and RSD loss are employed, we observe that RPR gradually increases incrementally with an increase in $\lambda_{rel}$. This suggests that the student model partially alleviates the dimensional collapse which is guided by a teacher model that has been trained with an entropy regularizer (SSCD). Subsequently, in the second section, where HN loss is incorporated, the student model finally utilizes the full dimensions of the descriptors.
Table 7. Ablation study to evaluate the impact of the loss ratio. All experiments are performed with a descriptor size of 256.
@@
We next conduct a step-by-step experiment with DINO as the teacher model (Table 8). Our approach reduces the descriptor size by a factor of 12 but produces a performance comparable to that of the teacher model. As illustrated in Discussion 5.1, dimensional collapse occurs when HN loss is not employed. Applying HN loss thus successfully improves performance and prevents dimensional collapse.
Table 8. Ablation study of using DINO as teacher model.
Method Network Size RPR $\mu$AP $\mu$APSN
DINO (Teacher) ViT-B/8 1536 0.99 32.6 54.4
@@
RSD — MoCo-v2 EN-B0 128 0.57 20.7 40.0
RSD — MoCo-v2 — HN (RDCD) EN-B0 128 1.0 32.1 52.1
6. Conclusion
@@
Appendix A Different Architectures
To show the generality of our method, we adopt one lightweight model each from the most commonly used architectures in computer vision, CNN and ViT, conducting experiments with MobileNet-V3 and FastViT-T12. Our results are shown in Table 9, Table 10, Table 11.
Table 9. Performance of DISC2021 with different student architectures. * means our implementation.
Method Network Size $\mu$AP $\mu$APSN
@@ Appendix A Different Architectures
Table 10. Performance of CD10K(Copydays+ 10k distractors) with different student architectures. * means our implementation.
Method Network Size mAP $\mu$AP
@@
Table 11. Performance of NDEC with different student architectures. * means our implementation.
Method Network Size $\mu$AP $\mu$APSN
@@
Appendix B Comparison with Barlow-Distill
We introduce the RDCD method using the MoCo framework’s negative sample queue in our distillation process. Alternatively, there are self-supervised learning (SSL) methods that do not employ negative samples and instead focus solely on positive samples. Inspired by the Barlow Twins approach, we have integrated the capability of distillation in our process by generating positive pairs within a batch through augmentation and optimizing the cross-correlation matrix between these pairs. This method, we call Barlow-Distill, ensures that each feature operates as independently as possible from the others, thereby minimizing feature redundancy. This minimization ensures that each feature retains unique information, leading to a richer data representation.
The Barlow-Distill approach distinctly avoids the use of negative samples, with all learning occurring through the adjustment of the correlation between positive pairs. This method is particularly effective when generating different transformations of the same image through augmentation techniques. Consequently, the model enhances the representation of the same object or scene in various forms. Barlow-Distill generalizes well by minimizing local features in two distinct augmented images. This process effectively captures higher levels of abstract information, ensuring that the latent representation retains sufficient information to learn useful representations for various downstream tasks. However, this concentration on high-level features may result in the loss of crucial low-level details, which can degrade performance. This loss is particularly problematic in image copy detection tasks, where missing fine-grained visual details can lead to degraded performance.
In contrast, SSL methods utilizing contrastive learning could be more suitable for tasks such as image copy detection. This method, by clearly differentiating between positive and negative samples, enables the model to more effectively recognize and learn local detail and fine-grained visual differences. Contrastive learning enhances the model’s ability to distinguish between similar images, thus improving accuracy and efficiency in the image copy detection task. This enhancement is vital for capturing important local features in images, a crucial ability for image copy detection. According to our research results presented in Table 12, the RDCD model shows enhanced performance compared to Barlow-Distill.
Table 12. Comparision with RDCD and Barlow-Distill
Method Network Size $\mu$AP $\mu$APSN
Barlow-Distill EN-B0 64 44.1 48.3
Barlow-Distill EN-B0 128 51.9 59.0
Barlow-Distill EN-B0 256 54.8 64.4
RDCD EN-B0 64 43.9 53.5
RDCD EN-B0 128 50.0 61.1
RDCD EN-B0 256 52.7 65.7
Appendix C Intermediate Descriptor
Table 13. Comparison of descriptor size of 512. We do not use Hard Negative loss for SSCD method as they use Koleo regularizer.
@@ Appendix C Intermediate Descriptor
Appendix D Qualitative Examples
@@
Figure 6. Example of top-2 retrieval results from the DISC2021 dataset.
Appendix E PCA
The dimensionality of the descriptor plays a crucial role in optimizing the trade-off between matching time and accuracy during the matching phase of ICD. Previous approaches have employed PCA and PCA whitening to create compact descriptors. The post-processing technique of PCA whitening is known for its efficacy in making the descriptor distribution more uniform (Pizzi et al., 2022). Our RDCD method, trained via deep learning, exhibits superior performance compared to the baseline, which utilizes large-sized descriptors that have undergone PCA whitening. These results are detailed in Table  14.
Table 14. The following table demonstrates that when dimension reduction is applied to the descriptor through PCA whitening, our approach outperforms the existing baseline in terms of $\mu$AP performance. We do not apply Mixup. * means our implementation.
Method Network Descriptor Size $\mu$AP $\mu$APSN $\mu$APSN256 $\mu$APSN128 $\mu$APSN64
DINO ViT-B/8 1536 32.6 54.4 47.7 42.3 32.7
DINO ViT-B/16 1536 32.2 53.8 47.4 42.0 32.0
SSCD RN-50 512 43.6 72.5 66.0 56.3 38.3
SSCD ResNext-101 1024 63.7 75.3 65.5 54.0 34.8
using lightweight
SSCD EN-B0 64 38.2 48.5 48.5 48.5 48.5
SSCD EN-B0 128 43.5 56.2 56.2 56.2 44.4
SSCD EN-B0 256 46.0 59.8 59.8 53.8 42.3
SSCD EN-B0 512 43.6 61.1 58.0 52.0 41.9
Ours
RDCD EN-B0 64 43.9 53.5 53.5 53.5 53.5
RDCD EN-B0 128 50.0 61.1 61.1 61.1 53.1
RDCD EN-B0 256 52.7 65.7 65.7 61.5 52.2

### v5

@@ 5.3 Ablation Study
Effectiveness of Each Loss Term. To further assess the effectiveness of RDCD, we investigate the impact of different loss components on its performance, specifically, RSD loss and HN loss as shown in Table 4. In this section, we present an additional metric, the rank preserving ratio (RPR), which is the rank divided by the descriptor dimension. A low RPR indicates dimensional collapse, and RPR value of 1 indicates that the descriptor makes full use of all dimensions. All models are trained on the DISC2021 dataset with 100 epochs and SSCD-RN50-w/mixup is employed as the teacher model. Our result shows that with the addition of RSD loss and HN loss, the performance improves substantially, which indicates that RSD and HN loss can benefit the performance of RDCD.
@@
Table 4: Ablation study on the loss terms in RDCD. Descriptor sizes of 64, 128, and 256 were examined, and a consistent MLP configuration was maintained across all experiments for a fair comparison.
@@
Influence of $\lambda_{rel}$ and $\lambda_{hn}$. $\lambda_{rel}$ and $\lambda_{hn}$ are the weights of RSD loss and HN loss, respectively. First, we analyze the influence of contrastive loss and RSD loss, as shown in Table 5. As shown in the third row, a contrastive loss of 1 and RSD loss of 10 achieves the best performance, which indicates that knowledge distillation is essential for the model with better performance. The HN loss is also varied with the contrastive loss and RSD loss fixed at 1 and 10, respectively, showing that the best performance occurs when the HN loss is 5. However, for larger values, the performance drops to 0, (i.e. collapses) because the network focuses more on HN loss rather than contrastive and distillation loss. We further calculate RPR of each experiment. In the first section, where only contrastive loss and RSD loss are employed, we observe that RPR gradually increases incrementally with an increase in $\lambda_{rel}$. This suggests that the student model partially alleviates the dimensional collapse which is guided by a teacher model that has been trained with an entropy regularizer. Subsequently, in the second section, where HN loss is incorporated, the student model finally utilizes the full dimensions of the descriptors.
Table 5: Ablation study to evaluate the impact of the loss ratio. All experiments are performed with a descriptor size of 256.
@@
We next conduct a step-by-step experiment with DINO as the teacher model, as shown in Table 6. Our approach reduces the descriptor size by a factor of 12 but produces a performance comparable to that of the teacher model. As illustrated in 5.1, dimensional collapse occurs when HN loss is not employed. Applying HN loss thus successfully improves performance and prevents dimensional collapse.
Table 6: Ablation study of using DINO as teacher model.
Method Network $dim$ RPR $\mu$AP $\mu$APSN
DINO [5] (Teacher) ViT-B/8 1536 0.99 32.6 54.4
@@
RSD |||| MoCo-v2 EN-B0 128 0.57 20.7 40.0
RSD |||| MoCo-v2 |||| HN (RDCD) EN-B0 128 1.0 32.1 52.1
6 Conclusion
@@
Acknowledgements
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (RS-2023-00224740, Development of technology to prevent and track the distribution of illegally filmed content).
@@ Acknowledgements
Appendix A Effect of Different Architectures
A.1 Different Student
To show the generality of RDCD, we adopt one lightweight model each from the most commonly used architectures in computer vision, CNN and ViT, conducting experiments with MobileNet-V3 and FastViT-T12. Our results are shown in Table 7, Table 8, Table 9. Specifically, using FastViT under the same conditions resulted in a $\mu$AP of 67.4, demonstrating the research potential for ViT student.
Table 7: Performance of DISC2021 with different student architectures. * means our implementation.
Method Network $dim$ $\mu$AP $\mu$APSN
@@ A.1 Different Student
Table 8: Performance of CD10K(Copydays+ 10k distractors) with different student architectures. * means our implementation.
Method Network $dim$ mAP $\mu$AP
@@
Table 9: Performance of NDEC with different student architectures. * means our implementation.
Method Network $dim$ $\mu$AP $\mu$APSN
@@
A.2 Different Teacher
We demonstrate the effectiveness of our proposed approach in scenarios where the teacher networks belong to different architectural families (ViT or CNN), with student models also selected from these two distinct architectures. For the ViT, we use DINO as the teacher, and for the CNNs, we use SSCD-RN50 and SSCD-ResNeXt-101. All models are trained on SSCD-RN50 hyperparameter settings. With DINO as the teacher model, we follow the ICD settings presented in [5] for the DINO baseline. Consequently, we distill a descriptor size of 1536 from the teacher (concatenation of 768 class tokens and 768 GeM pooled patch tokens). In the case of ResNeXt-101 teacher model, we distill the final target descriptor with a size of 1024. Table 10 shows that our method is highly effective for the different architectures, highlighting its adaptability and robustness.
Table 10: DISC2021 evaluation with different architectures. All experiments are performed with a descriptor size of 128.
T S EN-B0 FastViT-T12 Mob-V3
$\mu$AP $\mu$APSN $\mu$AP $\mu$APSN $\mu$AP $\mu$APSN
DINO-ViT-B/8 32.1 52.1 28.1 52.2 39.6 54.7
SSCD-RN-50 50.0 61.1 48.5 61.1 47.3 58.8
SSCD-ResNeXt-101 44.9 59.9 42.8 57.4 41.1 56.6
Table 11: The following table demonstrates that when dimension reduction is applied to the descriptor through PCA whitening, our approach outperforms the existing baseline in terms of $\mu$AP performance. We do not apply Mixup. * means our implementation.
Method Network $dim$ $\mu$AP $\mu$APSN $\mu$APSN256 $\mu$APSN128 $\mu$APSN64
DINO ViT-B/8 1536 32.6 54.4 47.7 42.3 32.7
DINO ViT-B/16 1536 32.2 53.8 47.4 42.0 32.0
SSCD RN-50 512 43.6 72.5 66.0 56.3 38.3
SSCD ResNext-101 1024 63.7 75.3 65.5 54.0 34.8
using lightweight
SSCD EN-B0 64 38.2 48.5 48.5 48.5 48.5
SSCD EN-B0 128 43.5 56.2 56.2 56.2 44.4
SSCD EN-B0 256 46.0 59.8 59.8 53.8 42.3
SSCD EN-B0 512 43.6 61.1 58.0 52.0 41.9
Ours
RDCD EN-B0 64 43.9 53.5 53.5 53.5 53.5
RDCD EN-B0 128 50.0 61.1 61.1 61.1 53.1
RDCD EN-B0 256 52.7 65.7 65.7 61.5 52.2
Table 12: Comparison of intermediate descriptor. We do not use HN loss for SSCD method as they use Koleo regularizer.
@@ A.2 Different Teacher
Appendix B Intermediate Descriptor
Appendix C PCA
The dimensionality of the descriptor plays a crucial role in optimizing the trade-off between matching time and accuracy during the matching phase of ICD. Previous approaches have employed PCA and PCA whitening to create compact descriptors. Following [31]The post-processing technique of PCA whitening is known for its efficacy in making the descriptor distribution more uniform. Our RDCD method, trained via deep learning, exhibits superior performance compared to the baseline, which utilizes large-sized descriptors that have undergone PCA whitening. These results are detailed in Table  11.
Table 13: Comparison of other distillation. All methods are trained with EfficientNet-B0 network. We do not use Hard Negative loss for SSCD method as they use Koleo regularizer.
128 descriptors
Hard Negative Loss No Yes
Method RSD FKD $\mu$AP $\mu$APSN $\mu$AP $\mu$APSN
SSCD 43.5 56.2 - -
SimCLR 46.2 58.7 48.4 59.9
47.4 59.6 49.6 61.4
47.5 59.6 50.8 61.5
MoCo-v2 40.3 57.5 44.8 59.1
47.1 60.7 50.0 61.1
46.2 59.3 49.6 61.0
256 descriptors
Hard Negative Loss No Yes
Method RSD FKD $\mu$AP $\mu$APSN $\mu$AP $\mu$APSN
SSCD 46.0 59.8 - -
SimCLR 51.7 64.0 51.4 64.5
52.2 65.6 52.7 65.8
52.4 65.5 49.9 64.6
MoCo-v2 41.3 56.6 46.1 60.1
47.2 61.0 52.9 65.7
47.3 60.8 53.2 65.9
Appendix D Comparison with Other Forms of Distillation
To further demonstrate the efficacy of RSD, we conduct a comparative analysis with FKD by 1) employing FKD alone, 2) employing RSD alone, and 3) integrating both FKD and RSD in Table 13. We also employ SimCLR-style contrastive learning. When FKD and RSD are combined, the best performance is observed in several cases, suggesting that combining both distillation methods can further improve performance. Nevertheless, RSD consistently outperforms FKD in all cases, with RSD alone achieving performance on par with the combined use of both distillation methods. This indicates that RSD is sufficiently robust, and FKD does not significantly alter the outcome. It confirms that RSD can operate alone to enhance the performance of compact descriptors.
Table 14: Ablation study on the computational cost during training. All experiments are conducted on a single A100 GPU using FP32 precision with EFF-B0 architecture and a descriptor size of 256. All metrics are calculated with a batch size of 256 except for FLOPs, which we measured with a batch size of 1.
Method Contrastive Learning Distillation #Params FLOPs Time per Epoch VRAM
(M) (G) (min) (G)
SSCD SimCLR X 4.3 0.828 58.46 49.7
- - O 29.3 8.732 40.06 28.1
- SimCLR O 29.4 9.16 74.34 54.9
RDCD MoCo-v2 O 34.2 9.9 51.25 29.8
Table 15: Comparison of computational cost during evaluation, highlighting our method’s high efficiency. We evaluate images with 288x288 size. FLOPs is calculated with a batch size of 1.
Method Architecture #Params FLOPs Throughput Search Time $dim$ $\mu$AP
(M) (G) (image/sec)
DINO ViT-B/8 85.204 133.699 87.76 37.5 1536 54.4
SSCD RN-50 24.6 8.265 903.58 15.518 512 72.5
RDCD EN-B0 4.8 0.829 1087.75 10.121 256 65.7
Appendix E Practicality of RDCD
We perform additional analysis of practicality on RDCD. To highlight the computational cost of each component, in Table 14 we conduct ablation experiments by removing contrastive learning and distillation. SSCD uses SimCLR without distillation. We found that our method requires 29.9 million more parameters than SSCD for training, primarily due to the large teacher network, which accounts for 24.0M parameters. Additionally, our method consumes a total of 9.9G of FLOPs: 8.3G for distillation and 1.6G for contrastive learning. Nevertheless, SSCD’s use of SimCLR leads to a proportional increase in VRAM usage with larger batch sizes. In our RDCD, we adopt MoCo-v2 which stores negative features in a queue, significantly reducing VRAM usage. By leveraging MoCo-v2, our method achieves 0.6 times lower VRAM usage compared to SSCD and also reduces training time. Furthermore, our final performance surpasses SSCD by 5.9.
Our RDCD incurs higher computational cost during training but shows significant efficiency during evaluation. Specifically, RDCD requires only 4.8M parameters during evaluation, which includes an 0.5M additional parameters from the EFF-B0 architecture. In copy detection, the evaluation involves two stages: model inference and database search. High-speed search is crucial when retrieving from a large-scale image database in real-world applications. We evaluate search time using the DISC2021 dataset, which consists of 1M database images and 50k query images. As shown in Table 15, RDCD achieves superior efficiency across all metrics compared to previous methods. DINO and SSCD require 133.699G, 8.265G FLOPs, respectively, due to their large architectures, while RDCD operates with 161 times lower FLOPs than DINO and 9.9 times lower FLOPs than SSCD during model inference.
Appendix F Qualitative Examples
@@
Figure 6: Example of top-2 retrieval results from the DISC2021 dataset.